OpenStudy (anonymous):

Researchers count 30 tagged zebras and 150 total zebras in an area of a game reserve. If researchers had previously tagged 80 total zebras at the game reserve, about how many total zebras live at the game reserve? Write a proportion to estimate the total population. (3 points)

OpenStudy (anonymous):

please help :(

OpenStudy (anonymous):

30/150 = 80/x where "x" is the total # of zebras. What you do now is solve for "x". Cross-multiplication is a good way to go. The ratio here was set up as showing the same proportion of sample to population.

OpenStudy (anonymous):

ok, so what you just wrote, on the top, it the equation i have to do? which is cross multiply?

OpenStudy (anonymous):

*is

OpenStudy (anonymous):

yes. Your problem asks only for an equation, not that you have to solve it, but it is good to do so. The reason "30 of 150" is a sample is because you look for that little word "area". 150 were in that "area" and area is used as a "sample" here.

OpenStudy (anonymous):

wow, ok, so when i do the math, i have to multipl 30, by x, and 150 by 80? :)

OpenStudy (anonymous):

*multiply

OpenStudy (anonymous):

Yes, that's the first step.

OpenStudy (anonymous):

and then what do i do?

OpenStudy (anonymous):

The second step is that you will take that 80 x 150 and divide that by 30 to get "x".

OpenStudy (anonymous):

Hint: as a check, both sides of that original equation will be equal to 1/5. That's the whole reasoning behind this whole problem. 1/5 in sample. 1/5 in population.

OpenStudy (anonymous):

ok, so 150, by 80 is....12,000?

OpenStudy (anonymous):

haha yeah

OpenStudy (anonymous):

was i right?

OpenStudy (anonymous):

80 times 150 = 12000 -> 12000 / 30 = 400 So, 400 in the whole reserve as an estimate.

OpenStudy (anonymous):

wow, your really good, thanks so much! :D this really helped

OpenStudy (anonymous):

uw! Good luck to you in all of your studies and thx for the recognition! @gabaloolo

OpenStudy (anonymous):

no probs, and also thank you :)